OpenStudy (anonymous):

(a^3)^4/(a^7)^6 simply

OpenStudy (anonymous):

start with \[\frac{a^{12}}{a^{42}}\] then cancel some

OpenStudy (anonymous):

a^3*4/a^7*6

OpenStudy (anonymous):

a^30

OpenStudy (anonymous):

should end up with \[\frac{1}{a^{42-12}}=\frac{1}{a^{30}}\]

OpenStudy (anonymous):

@hannalafave it is in the denominator, not the numerator

OpenStudy (anonymous):

it worked ++ thanks